# GOProteinGNN: Leveraging Protein Knowledge Graphs for Protein Representation Learning

### v1

@@ Abstract.
Proteins play a vital role in biological processes and are indispensable for living organisms. Accurate representation of proteins is crucial, especially in drug development. Recently, there has been a notable increase in interest in utilizing machine learning and deep learning techniques for unsupervised learning of protein representations. However, these approaches often focus solely on the amino acid sequence of proteins and lack factual knowledge about proteins and their interactions, thus limiting their performance. In this study, we present GOProteinGNN, a novel architecture that enhances protein language models by integrating protein knowledge graph information during the creation of amino acid level representations. Our approach allows for the integration of information at both the individual amino acid level and the entire protein level, enabling a comprehensive and effective learning process through graph-based learning. By doing so, we can capture complex relationships and dependencies between proteins and their functional annotations, resulting in more robust and contextually enriched protein representations. Unlike previous fusion methods, GOProteinGNN uniquely learns the entire protein knowledge graph during training, which allows it to capture broader relational nuances and dependencies beyond mere triplets as done in previous work. We perform a comprehensive evaluation on several downstream tasks demonstrating that GOProteinGNN consistently outperforms previous methods, showcasing its effectiveness and establishing it as a state-of-the-art solution for protein representation learning.
@@ 1. Introduction
Proteins play a vital role in biological processes and are central to understanding the intricate mechanisms of living organisms. In the pursuit of unraveling their complexity, the challenge of accurately representing proteins has gained considerable attention. Accurate representation holds paramount significance, especially for drug development (Liu et al., 2024).
@@
Amino acids serve as the fundamental building blocks of proteins. In recent years, there has been a surge of interest in employing machine learning and deep learning techniques to create unsupervised amino acid representations that can be used to represent a protein (Rives et al., 2019; Rao et al., 2021; Xiao et al., 2021). Prevailing approaches have predominantly focused on unraveling amino acid relationships using text-based methodologies, such as ProtBert (Elnaggar et al., 2022). The amino acid representations are then averaged to create the protein representation.
@@
Recently, numerous approaches have attempted to directly create a protein representation by encapsulating external information about the protein, rather than relying solely on the average of amino acid representations. Those methods, such as KeAP (Zhou et al., 2023), harness vast protein knowledge repositories. These knowledge graphs contain diverse information about interactions the protein participated in, the processes in which it is involved, etc. The factual knowledge is conveyed through text descriptions. Figure 3 provides an illustrative example of such a protein knowledge graph, showcasing the intricate connections between the protein, relations, and Gene Ontology (GO) terms. However, current state-of-the-art (SOTA) methods still grapple with the incorporation of this extensive information. Existing paradigms, at times, reduce the intricate protein knowledge graph to mere triplets (Protein, Relation, GO Term), neglecting the intricate tapestry of the knowledge graph (Zhou et al., 2023). These methods have overlooked the importance of capturing a holistic graph representation. For instance, although they may encode annotations, crucial information such as broader relational nuances and contextual dependencies within the graph may be lost unless the entire graph structure is considered. Moreover, these studies have predominantly concentrated on either amino acid level learning (Zhou et al., 2023; Elnaggar et al., 2022) or protein level learning (Zhang et al., 2022; Lam et al., 2023), with none addressing both simultaneously.
@@
In this work, we present GOProteinGNN, a novel architecture that integrates amino acid level and protein level representations within a single framework. Our method addresses existing limitations by comprehensively utilizing the entire protein knowledge graph structure during pretraining, ensuring the capture of intricate relationships. Specifically, GOProteinGNN leverages a novel Graph Neural Networks (GNN) Knowledge Injection (GKI) procedure that leverages the [CLS] token representation of the protein language model (PLM) during the encoder stage. GOProteinGNN interprets amino acid sequences as coherent “sentences” that describe proteins, learning embeddings for individual amino acids in its initial encoder layers. The GKI layer then assigns embeddings to protein nodes in the protein knowledge graph and incorporates knowledge into these embeddings by leveraging neighboring embeddings through a Relational Graph Convolution Network (RGCN) (Schlichtkrull et al., 2017). This process alternates between amino acid level embedding and graph representation, enhancing the overall representation by allowing new insights to spread across the amino acid sequence and ensuring a thorough encoding that incorporates both sequential and graph-enhanced information.
Comparison to Prior Art: Traditionally, most methods represented proteins solely based on their amino acid level interactions. Such methods (Rives et al., 2019; Elnaggar et al., 2022; Brandes et al., 2021) typically generate embeddings focused on the chemical characteristics of the proteins. For instance, these embeddings are capable of capturing the chemical interactions between protein pairs. In this work, by incorporating a knowledge graph into our approach, we are able to extend the encoding to include biological information, such as the specific biological processes in which the protein participates. In terms comparable to natural language processing, our methodology introduces a novel technique for integrating external sentence knowledge into word-embedding learning. During the training phase, we directly infuse a knowledge graph representation into the amino acid embeddings, leveraging the [CLS] token representation for this novel fusion. This enhancement significantly improves the representations, even for novel proteins that are not present in the knowledge graph. Other methods have concentrated on injecting knowledge graphs but have simplified the problem to only involve pairs of nodes (Zhou et al., 2023), often neglecting the broader structure of the knowledge graph. This simplification could adversely affect the quality of protein embeddings. For example, as illustrated in Figure 3, the KeAP method (Zhou et al., 2023), which focuses solely on node pairs within the knowledge graph, might not adequately capture the complex interactions involved in a protein’s role in diverse biological activities such as intracellular cAMP-activated cation channel activity, visual perception, and signal transduction pathways. This limitation stems from KeAP’s failure to account for a holistic view of the protein’s interactions within a biological context, potentially leading to a fragmented understanding of its functional significance, unlike our approach.
@@
We perform an empirical evaluation over numerous protein tasks, spanning a broad spectrum of biological domains, comparing to SOTA methods and achieving SOTA performance across various benchmarks. We present ablation tests to better study the performance of the algorithm.
@@
The contributions of this work are threefold: (1) We present a novel algorithm for protein representation learning that allows the integration of knowledge graph information into PLMs. This enables a comprehensive and effective learning process through graph-based learning. (2) We introduce a novel GNN Knowledge Injection (GKI) layer that enables protein knowledge information to propagate through the amino acid sequence representations. In addition, we leveraged the [CLS] token representation, to integrate the GKI with the PLM in a novel manner. (3) We present an extensive empirical evaluation of our work over numerous protein tasks establishing SOTA results. Also, we contribute our code to the community  [^1] GitHub repository: https://github.com/kalifadan/GOProteinGNN.
@@ 2.1. Proteins Representation Learning
Proteins are essential for living organisms and have vital roles in biological processes. Accurate representation of proteins is crucial, particularly in fields such as drug development. In recent years, there has been a surge of interest in applying machine learning and deep learning techniques to learn protein representations in an unsupervised manner. The popularity gain of Transformers (Vaswani et al., 2017) introduced many novel PLMs such as ESM (Rives et al., 2019), MSA Transformer (Rao et al., 2021), ProteinBERT (Brandes et al., 2021), ProteinLM (Xiao et al., 2021) and ProtTrans (Elnaggar et al., 2022). These models were trained on data from UniRef (Suzek et al., 2007) which contains hundreds of billions of amino acids.
A protein’s biological mechanism is determined by its three-dimensional (3D) native structure (Dill and MacCallum, 2012). AlphaFold (Jumper et al., 2021; Senior et al., 2020) is a recent model predicting the 3D structure of a protein. A common SOTA method that resembles the AlphaFold architecture but tuned for protein representation is ESM (Rives et al., 2019). Unlike AlphaFold, which primarily focuses on 3D structure prediction and relies on multiple sequence alignments and template structures, ESM requires only a single input sequence (i.e., protein) and has demonstrated similar performance across various tasks. Our comparative analysis focuses on ESM, emphasizing its utilization of single-sequence input and acknowledging its performance similarity to AlphaFold.
@@ 2.2. Knowledge-Enhanced Language Models
While PLMs traditionally learn protein representations from sequences, (Xu et al., 2023) proposed learning from biomedical text descriptions. Similarly, OntoProtein (Zhang et al., 2022) integrated external knowledge graphs into protein pre-training, focusing on Gene Ontology (GO) term and protein entity alignment. KeAP (Zhou et al., 2023) explored more granular knowledge graph interaction for improved protein decoding. OntoProtein’s contrastive learning and masked modeling were enriched by KeAP’s focus on the masked language modeling objective, achieving superior performance in various tasks. Otter-Knowledge (Lam et al., 2023) extended knowledge graph integration beyond proteins, incorporating diverse entities into a comprehensive knowledge graph.
The aforementioned models predominantly incorporate knowledge post-encoder, either via loss objectives or pre-training enrichment. This limitation hinders their ability to effectively incorporate external knowledge into protein embeddings during the encoder stage. Both OntoProtein (Zhang et al., 2022) and KeAP (Zhou et al., 2023), together with other existing models, have predominantly focused on either amino acid level learning or protein level learning. Also, they reduce intricate protein knowledge graph to mere triplets, neglecting the intricate tapestry of the knowledge graph. GOProteinGNN sets itself apart by uniquely encompassing both levels of representation within a single framework thus learning the entire protein knowledge graph and allowing it to capture broader relational nuances and dependencies beyond mere triplets.
@@ 3. Methods
(a) The GKI layer.
(b) The GOProteinGNN pre-training architecture.
Figure 2. The GOProteinGNN pre-training architecture. The model incorporates a protein’s knowledge graph with relations and GO terms. It uses encoder layers to process the amino acid sequence, producing amino acid representations and a protein embedding (which is represented by the [CLS] embedding). Then, the GKI refines the protein’s embedding with knowledge graph information using graph learning techniques, yielding a knowledge-enhanced protein representation that captures essential biological contexts and interactions. Subsequent encoder layers further enhance the amino acid representation, while the model performs masked language modeling (MLM) to restore masked amino acids.
3.1. GOProteinGNN

### v2

@@ Abstract.
Proteins play a vital role in biological processes and are indispensable for living organisms. Accurate representation of proteins is crucial, especially in drug development. Recently, there has been a notable increase in interest in utilizing machine learning and deep learning techniques for unsupervised learning of protein representations. However, these approaches often focus solely on the amino acid sequence of proteins and lack factual knowledge about proteins and their interactions, thus limiting their performance. In this study, we present GOProteinGNN, a novel architecture that enhances protein language models by integrating protein knowledge graph information during the creation of amino acid level representations. Our approach allows for the integration of information at both the individual amino acid level and the entire protein level, enabling a comprehensive and effective learning process through graph-based learning. By doing so, we can capture complex relationships and dependencies between proteins and their functional annotations, resulting in more robust and contextually enriched protein representations. Unlike previous methods, GOProteinGNN uniquely learns the entire protein knowledge graph during training, which allows it to capture broader relational nuances and dependencies beyond mere triplets as done in previous work. We perform a comprehensive evaluation on several downstream tasks demonstrating that GOProteinGNN consistently outperforms previous methods, showcasing its effectiveness and establishing it as a state-of-the-art solution for protein representation learning.
@@ 1. Introduction
Proteins play a vital role in biological processes and are central to understanding the mechanisms of living organisms. In the pursuit of unraveling their complexity, the challenge of accurately representing proteins has gained considerable attention. Accurate representation holds paramount significance, especially for drug development (Liu et al., 2024).
@@
Amino acids serve as the fundamental building blocks of proteins. In recent years, there has been a surge of interest in employing deep learning techniques to create unsupervised amino acid representations that can be used to represent a protein (Rives et al., 2019; Rao et al., 2021). Prevailing approaches have predominantly focused on unraveling amino acid relationships using text-based methodologies, such as ProtBert (Elnaggar et al., 2022). The amino acid representations are then averaged to create the protein representation.
@@
Recently, several approaches have attempted to directly create a protein representation by integrating external information about proteins, rather than relying solely on the amino acid representations. Those methods, such as KeAP (Zhou et al., 2023), harness vast protein knowledge graphs. These graphs contain diverse information about interactions the protein participated in, the processes in which it is involved, etc. The factual knowledge is conveyed through text descriptions. Figure 3 illustrates an example of such a protein knowledge graph, showcasing the intricate relations between proteins and Gene Ontology (GO) terms. However, current state-of-the-art (SOTA) methods still encounter challenges in effectively incorporating this extensive information.
Existing paradigms, at times, reduce the intricate protein knowledge graph to mere triplets (Protein, Relation, GO Term), neglecting the complex tapestry of the knowledge graph (Zhou et al., 2023). These methods have overlooked the importance of capturing a holistic graph representation. For instance, although they may encode annotations, crucial information such as broader relational nuances and contextual dependencies within the graph may be lost unless the entire graph structure is considered. The aforementioned models also incorporate knowledge post-encoder, either via loss objectives or pre-training enrichment (Zhou et al., 2023; Elnaggar et al., 2022). Language models often struggle to accurately embed knowledge, particularly long-tail knowledge (Kandpal et al., 2022). This limitation hinders their ability to effectively incorporate external knowledge into protein representations during the encoder stage. Moreover, these models have predominantly concentrated on either amino acid level learning (Zhou et al., 2023; Elnaggar et al., 2022) or protein level learning (Zhang et al., 2022; Lam et al., 2023), with none addressing both simultaneously.
In this work, we present GOProteinGNN, a novel architecture that integrates amino acid level and protein level representations within a single framework. Our method addresses existing limitations by comprehensively utilizing the entire protein knowledge graph structure during pre-training, ensuring the capture of intricate relationships. The main technical advancement of GOProteinGNN lies in its ability to capture broader relational nuances and dependencies between protein attributes, which existing models often overlook. In a specific context, when an attribute is connected to another attribute, determining how to embed these connections during protein representation learning is a complex challenge. Compared to GOProteinGNN, which addresses this limitation, KeAP (Zhou et al., 2023) and previous studies have failed to account for a holistic view of the protein’s interactions within a biological context, potentially leading to a fragmented understanding of its functional significance.
A key innovation of our approach is the integration of knowledge graph information into protein language models (PLMs), using our novel Graph Neural Networks (GNN) Knowledge Injection (GKI) mechanism, which leverages the [CLS] token representation during the encoder stage. This unique integration enables holistic learning of protein representations, as reflected in our SOTA results over extensive real-world bioinformatics tasks.
@@
Specifically, GOProteinGNN interprets amino acid sequences as coherent “sentences” that describe proteins, learning representations for individual amino acids in its initial encoder layers. The GKI layer then assigns representations to protein nodes in the protein knowledge graph and incorporates knowledge into these representations by leveraging neighboring representations through a Relational Graph Convolution Network (RGCN) (Schlichtkrull et al., 2017). These refined representations are processed by subsequent encoder layers, generating comprehensive amino acid representation. This process alternates between amino acid level representations and graph representation, enhancing the overall representation by allowing new insights to spread across the amino acid sequence and ensuring a thorough encoding that incorporates both sequential and graph-enhanced information.
@@
We perform an empirical evaluation over numerous protein tasks, spanning a broad spectrum of biological domains, comparing to superior methods and achieving SOTA performance across various benchmarks. We study the performance of our method through ablation tests.
@@
The contributions of this work are threefold: (1) We present a novel algorithm for protein representation learning that allows the integration of knowledge graph information into PLMs during the encoder stage. This enables a comprehensive and effective learning process through graph-based learning. (2) We introduce a novel GNN Knowledge Injection (GKI) mechanism that enables protein knowledge information to propagate through the amino acid sequence representations. We leverage the [CLS] token representation to integrate the GKI with the PLM in a novel manner. (3) We present an extensive empirical evaluation of our work over numerous real-world bioinformatics tasks establishing SOTA results. Also, we contribute our code to the community  [^1] https://github.com/kalifadan/GOProteinGNN.
@@
(a) The GKI layer.
(b) The GOProteinGNN architecture.
Figure 2. The GOProteinGNN pre-training architecture. The model incorporates a protein’s knowledge graph with relations and GO terms. It uses encoder layers to process the amino acid sequence, producing amino acid representations and a protein representation (which is represented by the [CLS] representation). Then, the GKI refines the protein’s representation with knowledge graph information using graph learning techniques, yielding a knowledge-enhanced protein representation that captures essential biological contexts and interactions. Subsequent encoder layers further enhance the amino acid representation, while the model performs masked language modeling (MLM) to restore masked amino acids.
@@ 2.1. Proteins Representation Learning
Proteins are essential for living organisms and have vital roles in biological processes. Accurate representation of proteins is crucial, particularly in fields such as drug development. In recent years, there has been a surge of interest in applying machine learning and deep learning techniques to learn protein representations in an unsupervised manner. The popularity gain of Transformers (Vaswani et al., 2017) introduced many novel PLMs such as ESM-1b (Rives et al., 2019), ESM-2 (Lin et al., 2023), MSA Transformer (Rao et al., 2021), ProteinBERT (Brandes et al., 2021), ProteinLM (Xiao et al., 2021) and ProtTrans (Elnaggar et al., 2022). These models were trained on data from UniRef (Suzek et al., 2007) which contains hundreds of billions of amino acids.
A protein’s biological mechanism is determined by its three-dimensional (3D) native structure (Dill and MacCallum, 2012). AlphaFold (Jumper et al., 2021; Senior et al., 2020) is a recent model predicting the 3D structure of a protein. A common SOTA method that resembles the AlphaFold architecture but tuned for protein representation is ESM (Rives et al., 2019; Lin et al., 2023). Unlike AlphaFold, which primarily focuses on 3D structure prediction and relies on multiple sequence alignments and template structures, ESM requires only a single input sequence (i.e., protein) and has demonstrated similar performance across various tasks. Our comparative analysis focuses on ESM, emphasizing its utilization of single-sequence input and acknowledging its performance similarity to AlphaFold.
@@ 2.2. Knowledge-Enhanced Language Models
While PLMs traditionally learn protein representations from sequences, ProtST (Xu et al., 2023) was trained on biomedical text descriptions. Similarly, OntoProtein (Zhang et al., 2022) integrated external knowledge graphs into protein pre-training, focusing on Gene Ontology (GO) term and protein entity alignment. KeAP (Zhou et al., 2023) explored more granular knowledge graph interaction for improved protein decoding. OntoProtein’s contrastive learning and masked modeling were enriched by KeAP’s focus on the masked language modeling objective, achieving superior performance in various tasks. Otter-Knowledge (Lam et al., 2023) extended knowledge graph integration beyond proteins, incorporating diverse entities into a comprehensive knowledge graph.
The aforementioned models also incorporate knowledge post-encoder, either via loss objectives or pre-training enrichment (Zhou et al., 2023; Elnaggar et al., 2022). Language models often struggle to accurately embed knowledge, particularly long-tail knowledge (Kandpal et al., 2022). This limitation hinders their ability to effectively incorporate external knowledge into protein representations during the encoder stage.
Both OntoProtein (Zhang et al., 2022) and KeAP (Zhou et al., 2023), together with other existing models, have predominantly focused on either amino acid level learning or protein level learning. Also, they reduce the intricate protein knowledge graph to mere triplets, neglecting the intricate tapestry of the knowledge graph. GOProteinGNN sets itself apart by uniquely encompassing both levels of representation within a single framework thus learning the entire protein knowledge graph and allowing it to capture broader relational nuances and dependencies beyond mere triplets.